# On the Wasserstein distance

### v2

@@ 5 A generalization: The Dobrushin-Shlosman Theorem
$\begin{split}\rho_{W}\left(\mu_{j}^{\beta}(\cdot\,|\underline{s}),\mu_{j}^{\beta}(\cdot\,|\underline{s}^{\prime})\right)=&\frac{1}{2}\left|\frac{\mathrm{e}^{4\beta J}}{Z_{\beta}(1,1,1,1)}-\frac{\mathrm{e}^{3\beta J}}{Z_{\beta}(1,1,1,2)}\right|\\ &+\frac{1}{2}\left|\frac{1}{Z_{\beta}(1,1,1,1)}-\frac{\mathrm{e}^{\beta J}}{Z_{\beta}(1,1,1,2)}\right|+\frac{1}{2}\left|\frac{1}{Z_{\beta}(1,1,1,1)}-\frac{1}{Z_{\beta}(1,1,1,2)}\right|,\end{split}$
@@
In dimension $d=3$, there are 23 combinations of boundary conditions to be considered. We simply show the Dobrushin lower bound in a graph analogous to Figure 3:
@@
The asymptotics is entirely analogous,
@@
6.3 Dobrushin-Shlosmann theorem and linear programming.
In order to apply the Dobrushin-Shlosmann theorem (Theorem 5.1) to some lattice spin models, we need a method, possibly computational, to estimate the Wasserstein distance for collections of spins, see, e.g., [25]. This problem is more generally known as linear programming, see, e.g., [33] and references therein.
@@
Improved lower bounds for $\beta_{c}J$ may be obtained by applying Dobrushin-Shlosman’ conditions (5.2)-(5.3) in Theorem 5.1. For illustration purposes, we consider the following cases: a square of $2\times 2$ spins and a square of $3\times 3$ spins for $d=2$, and a cube of $2\times 2\times 2$ spins for $d=3$. The linear programming technique was applied to compute the Wasserstein distance numerically for all relevant configurations of boundary conditions. Our results are summarised in the table below
@@
Each pair of boundary conditions (bc1 and bc2) differs by one single pair of corner boundary spins. For the calculation of the resulting Wasserstein distance, a weighting accounting for the number of spin changes is needed. In the square of $2\times 2$ spins configuration, the maximum Wasserstein distance equals $\frac{1}{2}$ (4 spins in $\Lambda_{0}$, 8 boundary spins), see (5.2)-(5.3). By computing the Wasserstein distance using the linear programming algorithm and then optimizing for $\beta J$, we obtain the lower bound $\beta_{c}J>0.30817$ for the critical inverse temperature. While the Wasserstein distance was computed for all possible boundary conditions, only a selection is given in the table below.
@@
The pair of boundary conditions only differs by one middle boundary spin. For the calculation of the resulting Wasserstein distance, a weighting accounting for the number of spin changes (corner spins vs middle spins) is needed. In the square of $3\times 3$ spins, the maximum Wasserstein distance equals $\frac{3}{4}$ (9 spins in $\Lambda_{0}$, 12 boundary spins), see (5.2)-(5.3). By computing the Wasserstein distance using the linear programming algorithm and then optimizing for $\beta J$, we get the lower bound $\beta_{c}J>0.33021$ for the critical inverse temperature. Due to the number of possible combinations, the computation of the unweighted Wasserstein distance was limited to a selection of relevant boundary conditions. The main cases are presented in the table below
@@
Noting that there are 12 possible changes, each corresponding to a sign flip of a different middle boundary spin and corner boundary spin, and each change may result in a different unweighted Wasserstein distance, the resulting Wasserstein distance is computed below for the 2 sets of boundary conditions from Table 3
@@
Each pair of boundary conditions (bc1 and bc2) only differs by one single triplet of corner boundary spins. For the calculation of the resulting Wasserstein distance, a weighting accounting for the number of spin changes is needed. In the cube of $2\times 2\times 2$ spins configuration, the maximum Wasserstein distance equals $\frac{1}{3}$ (8 spins in $\Lambda_{0}$, 24 boundary spins), see (5.2)-(5.3). By computing the Wasserstein distance using the linear programming algorithm and then optimizing for $\beta J$, we obtain the lower bound $\beta_{c}J>0.18727$ for the critical inverse temperature. Due to the number of possible combinations, the computation of the Wasserstein distance was limited to a selection of relevant boundary conditions. The main cases are presented in the table below
@@
Noting that there are 24 possible changes, each corresponding to a sign flip of a different corner boundary spin, and each change may result in a different unweighted Wasserstein distance, the resulting Wasserstein distance is computed below for the first set of boundary conditions from Table 5
@@
Then, provided that $\alpha>2d\|L\|$, there exists a unique limit Gibbs distribution at inverse temperature $\beta>0$ in $\mathcal{P}((\mathbb{R}^{n})^{\mathbb{Z}^{d}})$, associated with $H_{n}^{\textrm{cla}}$, with marginal distributions satisfying (3.11).
@@
Assume that $g:\mathbb{R}\rightarrow\mathbb{R}$ is a convex function. Then provided that $\alpha>0$, there exists a unique limit Gibbs distribution at inverse temperature $\beta>0$ in $\mathcal{P}(\mathbb{R}^{\mathbb{Z}^{d}})$, associated with $\tilde{H}_{1}^{\textrm{cla}}$, with marginal distributions satisfying (3.11).
@@
×$\int_{\mathbb{R}\times\mathbb{R}}|x-x^{\prime}|\sigma_{j;\underline{y},\underline{y}^{\prime}}^{\beta}(dx,dx^{\prime})=\sum_{m\in\mathbb{Z}}\int_{(z_{m-1},z_{m}]\times(z_{m-1}^{\prime},z_{m}^{\prime}]}|x-x^{\prime}|\frac{\mu^{\beta}_{j}(dx|\underline{y})\mu^{\beta}_{j}(dx^{\prime}|\underline{y}^{\prime})}{\mu^{\beta}_{j}\left((z_{m-1},z_{m}]|\underline{y}\right)}\\ \leq\sum_{m\in\mathbb{Z}}\left(\sup_{(z,z^{\prime})\in(z_{m-1},z_{m}]\times(z_{m-1}^{\prime},z_{m}^{\prime}]}|z-z^{\prime}|\right)\frac{\mu^{\beta}_{j}\left((z_{m-1},z_{m}]|\underline{y}\right)\mu^{\beta}_{j}\left((z_{m-1}^{\prime},z_{m}^{\prime}]|\underline{y}^{\prime}\right)}{\mu^{\beta}_{j}\left((z_{m-1},z_{m}]|\underline{y}\right)}.$ (7.10)
@@
Assume that $g_{r}:\mathbb{R}\rightarrow\mathbb{R}$, $r=1\ldots,n$ are convex functions. Then, provided that $\alpha>2d\|L\|$, there exists a unique limit Gibbs distribution at inverse temperature $\beta>0$ in $\mathcal{P}((\mathbb{R}^{n})^{\mathbb{Z}^{d}})$, associated with $\tilde{H}_{n}^{\textrm{cla}}$, with marginal distributions satisfying (3.11).
@@
Some translation-invariant Hamiltonians with infinite-range pair-interactions are considered in [50, 17, 11]. A typical pair-interaction potential is of the form (7.5). In [50, 17, 11], the self-interaction potentials diverge at least quadratically and the conditions set on the pair-interaction potentials assure the superstability of the Hamiltonians. For a wide class of boundary conditions, existence of ’superstable’ limit Gibbs distributions (see [69]) is proven in [50] (the state space is $\mathbb{R}^{n}$) and existence of tempered limit Gibbs distributions is proven in [17, 11] (the state space is $\mathbb{R}$). Conditions for uniqueness are also discussed in [50, 17, 11]. In [17], sufficient conditions are derived from Dobrushin’s uniqueness theorem using an explicit expression for the Wasserstein distance between probability measures on $\mathbb{R}$, see [17, Thm. 2.2].
@@
Let $\mathcal{X}$ be a completely regular Hausdorff space. Let $K\subset\mathcal{X}$ be a non-empty compact set and $O_{1},O_{2}\subset\mathcal{X}$ be two open sets such that $K\subset(O_{1}\cup O_{2})$. Let $(\mu_{n})_{n\in\mathbb{N}}\subset\mathcal{P}(\mathcal{X})$ be a sequence of probability measures converging weakly to $\mu\in\mathcal{P}(\mathcal{X})$. Then, given $10<\delta<$ 1 and $10<\epsilon<$ 1, there exist two open sets $V_{1}\subset O_{1}$ and $V_{2}\subset O_{2}$ such that $K\subset(V_{1}\cup V_{2})$, with $(K\setminus O_{2})\subset V_{1}$ and $(K\setminus V_{1})\subset V_{2}$, such that, for all $n$ large enough, all the following hold
@@
that we have $\sigma_{n}\in\Xi_{\mathcal{X}}(\mu,\mu_{n})$. Moreover, we have,
@@
9.2.1 The dual problem.
@@
$=\sup_{(\underline{x}\underline{y})\in\mathbb{R}^{n}\times\mathbb{R}^{m}}\left(\left\langle\underline{y},\underline{v}\right\rangle_{\mathbb{R}^{m}}-F\left(\underline{x},\underline{y}\right)\right)=F^{*}\left(\underline{0},\underline{v}\right).$square
@@
$\rho_{W}(\mu,\nu)=\sup_{f\in{\mathrm{Lip}}_{1}(\mathcal{X})}\int_{\mathcal{X}}f(x)\,\mu(dx)-\int_{\mathcal{X}}f(x)\,\nu(dx),$ (9.8)
@@
$G((\alpha,\beta)):=\begin{cases}0&\textrm{if $\alpha=(\mu-\nu)_{+}$ and $\beta=(\mu-\nu)-$;}\\ +\infty&\text{otherwise.}\end{cases}$ (9.10)
@@
$\tilde{f}(y):=\begin{cases}\sup_{x\in\mathcal{X}_{+}}\left(f(x)-\rho(x,y)\right)&,\textrm{if $f(y)>f(x_{0})+\rho(x_{0},y)$;}\\ f(y)&,\textrm{otherwise.}\end{cases}$
@@
Then, provided that $\alpha>0$, there exists a unique limit (Euclidean) Gibbs distribution at inverse temperature $\beta>0$ in $\mathcal{P}((L^{2}(\mathbb{T}_{\beta}))^{\mathbb{Z}^{d}})$, associated with $h^{\textrm{qua}}$, with marginal distributions satisfying (3.11).
@@
$Z_{j}^{\beta}(\underline{\eta}):=\int_{L^{2}(\mathbb{T}_{\beta})}\exp\left(-\frac{1}{2}\alpha_{d}\|\omega\|_{2}^{2}+\langle\omega,\eta\rangle_{\beta}\right)\mu_{0}^{\beta}(d\omega).$
@@
In view of (LABEL:Zfunc)-(9.17), by definition of Wiener measures, see, e.g., [16, Sec. 6.3.2], we have,

### v4

@@ 5 A generalization: The Dobrushin-Shlosman Theorem
$\begin{split}\rho_{W}\left(\mu_{j}^{\beta}(\cdot\,|\underline{s}),\mu_{j}^{\beta}(\cdot\,|\underline{s}^{\prime})\right)=&\frac{1}{2}\left|\frac{\mathrm{e}^{4\beta J}}{Z_{\beta}(1,1,1,1)}-\frac{\mathrm{e}^{3\beta J}}{Z_{\beta}(1,1,1,2)}\right|\\ &+\frac{1}{2}\left|\frac{1}{Z_{\beta}(1,1,1,1)}-\frac{\mathrm{e}^{\beta J}}{Z_{\beta}(1,1,1,2)}\right|+\frac{q-2}{2}\left|\frac{1}{Z_{\beta}(1,1,1,1)}-\frac{1}{Z_{\beta}(1,1,1,2)}\right|,\end{split}$
@@
In dimension $d=3$, there are 23 combinations of boundary conditions to be considered. We simply show the Dobrushin lower bound in Figure 4 below analogous to Figure 3.
@@
The asymptotics is entirely analogously,
@@
6.3 Dobrushin-Shlosman theorem and linear programming.
In order to apply the Dobrushin-Shlosman theorem (Theorem 5.1) to some lattice spin models, we need a method, possibly computational, to estimate the Wasserstein distance for collections of spins, see, e.g., [25]. This problem is more generally known as linear programming, see, e.g., [33] and references therein.
@@
Improved lower bounds for $\beta_{c}J$ may be obtained by applying Dobrushin-Shlosman’ conditions (5.2)-(5.3) in Theorem 5.1. For illustration purposes, we consider the following cases: a square of $2\times 2$ spins and a square of $3\times 3$ spins for $d=2$, and a cube of $2\times 2\times 2$ spins for $d=3$. The linear programming technique was applied to compute the Wasserstein distance numerically for all relevant configurations of boundary conditions. Our results are summarised in Table 1 below.
@@
Each pair of boundary conditions (bc1 and bc2) differs by one single pair of corner boundary spins. For the calculation of the resulting Wasserstein distance, a weighting accounting for the number of spin changes is needed. In the square of $2\times 2$ spins configuration, the maximum Wasserstein distance equals $\frac{1}{2}$ (4 spins in $\Lambda_{0}$, 8 boundary spins), see (5.2)-(5.3). By computing the Wasserstein distance using the linear programming algorithm and then optimizing for $\beta J$, we obtain the lower bound $\beta_{c}J>0.30817$ for the critical inverse temperature. While the Wasserstein distance was computed for all possible boundary conditions, only a selection is given in Table 2 below.
@@
The pair of boundary conditions only differs by one middle boundary spin. For the calculation of the resulting Wasserstein distance, a weighting accounting for the number of spin changes (corner spins vs middle spins) is needed. In the square of $3\times 3$ spins, the maximum Wasserstein distance equals $\frac{3}{4}$ (9 spins in $\Lambda_{0}$, 12 boundary spins), see (5.2)-(5.3). By computing the Wasserstein distance using the linear programming algorithm and then optimizing for $\beta J$, we get the lower bound $\beta_{c}J>0.33021$ for the critical inverse temperature. Due to the number of possible combinations, the computation of the unweighted Wasserstein distance was limited to a selection of relevant boundary conditions. The main cases are presented in Table 3 below.
@@
Noting that there are 12 possible changes, each corresponding to a sign flip of a different middle boundary spin and corner boundary spin, and each change may result in a different unweighted Wasserstein distance, the resulting Wasserstein distance is computed in Table 4 below for the 2 sets of boundary conditions from Table 3.
@@
Each pair of boundary conditions (bc1 and bc2) only differs by one single triplet of corner boundary spins. For the calculation of the resulting Wasserstein distance, a weighting accounting for the number of spin changes is needed. In the cube of $2\times 2\times 2$ spins configuration, the maximum Wasserstein distance equals $\frac{1}{3}$ (8 spins in $\Lambda_{0}$, 24 boundary spins), see (5.2)-(5.3). By computing the Wasserstein distance using the linear programming algorithm and then optimizing for $\beta J$, we obtain the lower bound $\beta_{c}J>0.18727$ for the critical inverse temperature. Due to the number of possible combinations, the computation of the Wasserstein distance was limited to a selection of relevant boundary conditions. The main cases are presented in Table 5 below.
@@
Noting that there are 24 possible changes, each corresponding to a sign flip of a different corner boundary spin, and each change may result in a different unweighted Wasserstein distance, the resulting Wasserstein distance is computed in Table 6 below for the first set of boundary conditions from Table 5.
@@
Then, provided that $\alpha>2d\|L\|$, there exists a unique limit Gibbs distribution at any inverse temperature $\beta>0$ in $\mathcal{P}((\mathbb{R}^{n})^{\mathbb{Z}^{d}})$, associated with $H_{n}^{\textrm{cla}}$, with marginal distributions satisfying (3.11).
@@
Assume that $g:\mathbb{R}\rightarrow\mathbb{R}$ is a convex function. Then provided that $\alpha>0$, there exists a unique limit Gibbs distribution at any inverse temperature $\beta>0$ in $\mathcal{P}(\mathbb{R}^{\mathbb{Z}^{d}})$, associated with $\tilde{H}_{1}^{\textrm{cla}}$, with marginal distributions satisfying (3.11).
@@
×$\begin{split}\int_{\mathbb{R}\times\mathbb{R}}|x-x^{\prime}|\sigma_{j;\underline{y},\underline{y}^{\prime}}^{\beta}(dx,dx^{\prime})&=\sum_{m\in\mathbb{Z}}\int_{(z_{m-1},z_{m}]\times(z_{m-1}^{\prime},z_{m}^{\prime}]}|x-x^{\prime}|\frac{\mu^{\beta}_{j}(dx|\underline{y})\mu^{\beta}_{j}(dx^{\prime}|\underline{y}^{\prime})}{\mu^{\beta}_{j}\left((z_{m-1},z_{m}]|\underline{y}\right)}\\ &\leq\sum_{m\in\mathbb{Z}}\left(\sup_{(z,z^{\prime})\in(z_{m-1},z_{m}]\times(z_{m-1}^{\prime},z_{m}^{\prime}]}|z-z^{\prime}|\right)\mu^{\beta}_{j}\left((z_{m-1}^{\prime},z_{m}^{\prime}]|\underline{y}^{\prime}\right).\end{split}$ (7.10)
@@
Assume that $g_{r}:\mathbb{R}\rightarrow\mathbb{R}$, $r=1\ldots,n$ are convex functions. Then, provided that $\alpha>2d\|L\|$, there exists a unique limit Gibbs distribution at any inverse temperature $\beta>0$ in $\mathcal{P}((\mathbb{R}^{n})^{\mathbb{Z}^{d}})$, associated with $\tilde{H}_{n}^{\textrm{cla}}$, with marginal distributions satisfying (3.11).
@@
Some translation-invariant Hamiltonians with infinite-range pair-interactions are considered in [50, 17, 11]. A typical pair-interaction potential is of the form (7.5). In [50, 17, 11], the self-interaction potentials diverge at least quadratically and the conditions set on the pair-interaction potentials assure the superstability of the Hamiltonians. For a wide class of boundary conditions, existence of ’superstable’ limit Gibbs distributions (see [68, 69]) is proven in [50] (the state space is $\mathbb{R}^{n}$) and existence of tempered limit Gibbs distributions is proven in [17, 11] (the state space is $\mathbb{R}$). Conditions for uniqueness are also discussed in [50, 17, 11]. In [17], sufficient conditions are derived from Dobrushin’s uniqueness theorem using an explicit expression for the Wasserstein distance between probability measures on $\mathbb{R}$, see [76, Theorem] and [17, Thm. 2.2].
@@
Let $\mathcal{X}$ be a completely regular Hausdorff space. Let $K\subset\mathcal{X}$ be a non-empty compact set and $O_{1},O_{2}\subset\mathcal{X}$ be two open sets such that $K\subset(O_{1}\cup O_{2})$. Let $(\mu_{n})_{n\in\mathbb{N}}\subset\mathcal{P}(\mathcal{X})$ be a sequence of probability measures converging weakly to $\mu\in\mathcal{P}(\mathcal{X})$. Then, given $10<\delta<$ 1 and $10<\epsilon<$ 1, there exist two open sets $V_{1}\subset O_{1}$ and $V_{2}\subset O_{2}$ such that $K\subset(V_{1}\cup V_{2})$, with $(K\setminus O_{2})\subset V_{1}$ and $(K\setminus V_{1})\subset V_{2}$, and such that, for all $n$ large enough, all the following hold
@@
that we have $\sigma_{n}\in\Xi_{\mathcal{X}}(\mu,\mu_{n})$. Moreover,
@@
9 Acknowledgement.
The authors are grateful to the referees, whose remarks led us to make significant improvements in our original manuscript.
@@
10.2.1 The dual problem in linear programming.
@@
$=\sup_{(\underline{x},\underline{y})\in\mathbb{R}^{n}\times\mathbb{R}^{m}}\left(\left\langle\underline{y},\underline{v}\right\rangle_{\mathbb{R}^{m}}-F\left(\underline{x},\underline{y}\right)\right)=F^{*}\left(\underline{0},\underline{v}\right).$square
@@
$\rho_{W}(\mu,\nu)=\sup_{f\in{\mathrm{Lip}}_{1}(\mathcal{X})}\left(\int_{\mathcal{X}}f(x)\,\mu(dx)-\int_{\mathcal{X}}f(x)\,\nu(dx)\right),$ (10.8)
@@
$G((\alpha,\beta)):=\begin{cases}0&\textrm{if $\alpha=(\mu-\nu)_{+}$ and $\beta=(\mu-\nu)_{-}$;}\\ +\infty&\text{otherwise.}\end{cases}$ (10.10)
@@
$\tilde{f}(y):=\begin{cases}\sup_{x\in\mathcal{X}_{+}}\left(f(x)-\rho(x,y)\right)&,\textrm{if $f(y)>f(x_{0})+\rho(x_{0},y)$;}\\ f(y)&,\textrm{otherwise.}\end{cases}$ (10.13)
@@
Then, provided that $\alpha>0$, there exists a unique limit (Euclidean) Gibbs distribution at any inverse temperature $\beta>0$ in $\mathcal{P}((L^{2}(\mathbb{T}_{\beta}))^{\mathbb{Z}^{d}})$, associated with $h^{\textrm{qua}}$, with marginal distributions satisfying (3.11).
@@
$Z_{j}^{\beta}(\underline{\eta}):=\int_{L^{2}(\mathbb{T}_{\beta})}\exp\left(-\frac{1}{2}\alpha_{d}\|\omega\|_{2}^{2}+\langle\omega,\eta\rangle_{\beta}\right)\mu_{0}^{\beta}(d\omega).$ (10.18)
@@
In view of (10.18)-(10.19), by definition of Wiener measures, see, e.g., [16, Sec. 6.3.2], we have,